# DESY 25-127, ZU-TH 59/25, MPP-2025-182, RISC Report Series 25–08 The three-loop single-mass heavy-flavor corrections to the structure functions $F_{2}(x,Q^{2})$ and $g_{1}(x,Q^{2})$

### v1

@@ Abstract
We report quantitative results on the single-mass heavy-flavor contributions up to three-loop order to the unpolarized structure function $F_{2}(x,Q^{2})$ and the polarized structure function $g_{1}(x,Q^{2})$ for the first time. These results are relevant for precision QCD analyses of the World deep-inelastic data and the data taken at future colliders, such as the Electron–Ion Collider, in order to measure the strong coupling constant $\alpha_{s}(M_{Z}^{2})$, and the twist-2 parton distribution functions consistently at next-to-next-to-leading order.
pacs:
12.38.Bx,12.38.Qk,14.65.Dw,14.65.Fy
I Introduction
@@
With the final results on the next-to-next-to-leading order (NNLO) single-mass corrections to the heavy-flavor Wilson coefficients in the region $Q^{2}\gg m_{Q}^{2}$, with $m_{Q}$ the heavy quark mass, first predictions on the structure functions $F_{2}$ and $g_{1}$ can be made at this order. The massive Wilson coefficients can be represented by Mellin convolutions of the massive operator matrix elements (OMEs)
$(\Delta)A_{ij}=\delta_{ij}+\sum_{k=1}^{\infty}a_{s}^{k}(\Delta)A_{ij}^{(k)}$ (1)
and pieces of the massless Wilson coefficients, cf. [9, 10]. The massive OMEs have been calculated to three-loop order in the upolarized [11, 12, 13, 14, 15, 16] and polarized cases, [12, 18, 17, 15, 16] [^1] The one– and two–loop corrections were calculated in Refs.  [19, 20, 21, 22, 23, 24] and [25, 26, 9, 27, 28, 29, 30, 31].
@@
Beyond the massive OMEs contributing to the massive Wilson coefficients, there are also others, $(\Delta)A_{gq,Q}$ and $(\Delta)A_{gg,Q}$, playing a role in the variable flavor number scheme [32, 14, 33, 35, 36, 34],
@@
It has been shown in Ref. [9] that for the structure function $F_{2}(x,Q^{2})$ the heavy-flavor corrections in the asymptotic region $Q^{2}\gg m_{Q}^{2}$ agree with the complete result at the level of $O(1\%)$ if $Q^{2}/m_{Q}^{2}>10$ at two-loop order. This cut also safely removes the higher-twist effects, cf. Refs. [37, 38, 39].
The structure function $F_{2}(x,Q^{2})$ consists of the massless part, $F_{2}^{\rm light}(x,Q^{2})$ for the light flavors, and the massive part, $F_{2}^{\rm heavy}(x,Q^{2})$, where the Wilson coefficients have either a charm or a bottom quark contribution.
@@
In our illustrations, we work in the fixed flavor number scheme for $N_{F}=3$ light flavors with charges $e_{k}$. The formal structures are the same for the structure functions $F_{2}(x,Q^{2})$ and $g_{1}(x,Q^{2})$, and are obtained by substituting the unpolarized building blocks with the polarized ones. For $g_{1}(x,Q^{2})$ the prefactor $x$ in Eqs. (3,I) has to be replaced by $1/2$.
The light flavor part of $F_{2}$ is given by [40, 41]
$F_{2}^{\rm light}(x,Q^{2})$ $=$ $x\sum_{k=0}^{N_{F}}e_{k}^{2}\Biggl\{\frac{1}{N_{F}}\Biggl[C_{2,q}^{\rm S}\left(x,\frac{Q^{2}}{\mu^{2}}\right)\otimes\Sigma(x,\mu^{2})$ (3)
$+C_{2,g}^{\rm S}\left(x,\frac{Q^{2}}{\mu^{2}}\right)\otimes G(x,\mu^{2})\Biggr]$
@@
$\Delta_{k}(x,\mu^{2})$ $=$ $f_{k}(x,\mu^{2})+\bar{f}_{k}(x,\mu^{2})-\frac{1}{N_{F}}\Sigma(x,\mu^{2}),$ (4)
$\Sigma(x,\mu^{2})$ $=$ $\sum_{k=1}^{N_{F}}[f_{k}(x,\mu^{2})+\bar{f}_{k}(x,\mu^{2})],$ (5)
with $f_{k,(\bar{k})}$ the quark and antiquark densities, and $\mu$ denotes the factorization scale. The unpolarized massless Wilson coefficients $C_{2,i}^{j}$ were calculated to three-loop order in Refs. [40, 41] and the polarized ones $\Delta C_{1,i}^{j}$ in Ref. [41]. The heavy-flavor part is given by, cf. Ref. [10],
$F_{2}^{Q}(x,Q^{2})=$
$x\Biggl\{\sum_{k=0}^{N_{F}}e_{k}^{2}\Biggl[L_{2,q}^{\rm NS}(x)\otimes(f_{k}(x,\mu^{2})+f_{\bar{k}}(x,\mu^{2}))$
$+\frac{1}{N_{F}}\left({L}_{2,q}^{\rm PS}(x)\otimes\Sigma(x,\mu^{2})+{L}_{2,g}^{\rm S}(x)\otimes G(x,\mu^{2})\right)\Biggr]$
@@
It contains both flavor non-singlet and singlet contributions [^2] For the study of pure non-singlet structure functions, see Ref. [42]. There the heavy flavor corrections are small. Here we have suppressed the dependence on $Q^{2},m_{Q}^{2}$ and $\mu^{2}$ in the heavy-flavor Wilson coefficients $H_{i}$ and $L_{i}$, and $\otimes$ denotes the Mellin convolution [^3] For the evaluation of the convolution integrals we use the integrator of Ref. [43].
$A(x)\otimes B(x)=\int_{0}^{1}dx_{1}\int_{0}^{1}dx_{2}\delta(x-x_{1}x_{2})A(x_{1})B(x_{2}).$
In Mellin space, the massive Wilson coefficients are polynomials of the expansion coefficients of the massive OMEs $(\Delta)A_{ij}$ and expansion coefficients of the massless Wilson coefficients, see Refs. [10, 14]. The massless Wilson coefficients to three-loop order can all be expressed in terms of either harmonic sums [44, 45] in Mellin space or harmonic polylogarithms [46] in $x$-space. For the massive Wilson coefficients, this only applies to the $N_{F}$-contributions [11] and to $(\Delta)L_{q}^{\rm NS},(\Delta)L_{q}^{\rm PS}$ and $(\Delta)L_{g}^{\rm S}$. The pure singlet Wilson coefficient $(\Delta)H_{q}^{\rm PS}$ depends on generalized harmonic sums [47, 48] in Mellin space and can be cast into harmonic polylogarithms with changed argument in $x$-space. The gluonic Wilson coefficients $(\Delta)H_{g}^{\rm S}$ receive also finite binomial sum [49] and higher transcendental function contributions in Mellin space. They contain square-root valued iterated integrals [49] and $_{2}F_{1}-$solutions [50] in $x$-space. [^4] For surveys on the different mathematical structures and the used computer algebraic methods see Refs. [51, 52]. Since we calculate the heavy-flavor corrections to inclusive structure functions, there are also contributions with massless final states, but virtual heavy-flavor corrections.
@@
II The Structure Function $F_{2}(x,Q^{2})$
In the unpolarized case, we use the parton distribution functions (PDFs) of Ref. [7] for $N_{F}=3$. For the values of $\alpha_{s}$ in the Wilson coefficients, we use flavor matching with the values $\alpha_{s}(25~\rm GeV^{2})=0.2020,\alpha_{s}(10^{2}~\rm GeV^{2})=0.1706$, and $\alpha_{s}(10^{3}~\rm GeV^{2})=0.1359$, corresponding to $\alpha_{s}(M_{Z}^{2})=0.1147$, consistent with the PDF fits. The massive Wilson coefficients were calculated referring to on-shell heavy quark masses. We use the following values
$m_{c}=1.59~\rm GeV,~~~~~{\it m_{b}}=4.78~\rm GeV,$ (8)
@@ II The Structure Function $F_{2}(x,Q^{2})$
In Figure 1, we illustrate the prediction for the total structure function $F_{2}^{\rm total}(x,Q^{2})$, including the single-mass charm and bottom quark contributions. In the region $x\sim 10^{-5}$, it rises from $Q^{2}=25~\rm GeV^{2}$ to $Q^{2}=1000~\rm GeV^{2}$ from values of $\sim 3.5$ to 10.5.
Figure 2: The different heavy-flavor contributions due to charm quarks to the structure function $F_{2}(x,Q^{2})$ by the Wilson coefficients $H_{g}^{\rm S},H_{q}^{\rm PS},L_{q}^{\rm NS},L_{g}^{\rm S}$ and $L_{g}^{\rm PS}$ at $Q^{2}=100~\rm GeV^{2}$ at different cumulative orders in $a_{s}$.
@@
Around $x=0.1$ there is a turning point above which the structure function takes lower values for larger values of $Q^{2}$.
Figure 2 shows the contributions of the different heavy-flavor Wilson coefficients to $F_{2}^{c}$ at $Q^{2}=100~\rm GeV^{2}$. The different lines illustrate also the contributions from $O(a_{s})$ to $O(a_{s}^{3})$ to the structure function. The largest contribution is due to the gluonic Wilson coefficient $H_{g}^{\rm S}$ which was scaled down by a factor of 20 for better readability. The corrections are positive and grow with the order in $a_{s}$. It is followed in size by a negative contribution from the pure-singlet Wilson coefficient $H_{q}^{\rm PS}$, which starts at $O(a_{s}^{2})$. In the large-$x$ region, the non-singlet Wilson coefficient $L_{q}^{\rm NS}$ yields the dominant part. However, the corrections compared to $F_{2}^{c,b}$ stay rather small. The corrections can become negative since virtual effects of heavy quarks with massless final states are contained in $F_{2}^{c}$ and $F_{2}^{b}$. Also the contributions of $L_{g}^{\rm S}$ and $L_{g}^{\rm PS}$ are of importance at the level of accuracies of $O(1\%)$. $L_{g}^{\rm S}$ is tiny at $O(a_{s}^{2})$, but receives a relative large correction at $O(a_{s}^{3})$. A part of Wilson coefficients emerging at higher order in the coupling constant $a_{s}$ can be negative as long as the unpolarized structure function remains positive.
In Figure 3 the heavy-flavor corrections to $F_{2}(x,Q^{2})$ are shown at NNLO. At $x=10^{-5}$, they grow from 1 to $\sim 4$ for $Q^{2}=25~\rm GeV^{2}$ to $Q^{2}=1000~\rm GeV^{2}$. In the large-$x$ region the corrections are negative.
@@
at NNLO. In the small-$x$ region of $x\sim 10^{-5}$, the fractions of the heavy-flavor corrections evolve from $Q^{2}=25~\rm GeV^{2}$ to $Q^{2}=1000~\rm GeV^{2}$ from 26 % to 41 %. At $x\sim 0.2$, the heavy-flavor corrections become zero, which is due to a combination of real and virtual corrections at NLO and NNLO. In the very-large-$x$ region, the heavy-flavor contributions at NNLO become negative and as large as $10\%$, but the structure function $F_{2}$ itself remains non-negative. It is clearly seen that the scaling violations of the massive contributions are rather different compared to those in the massless case up to high values of $Q^{2}$.

### v2

# DESY 25-127 arXiv:2509.16124 [hep-ph] ZU-TH 59/25 September 2025

@@ Abstract
We present quantitative results on the single-mass heavy-flavor contributions in the region of large virtualities $Q^{2}$ up to three-loop order to the unpolarized structure function $F_{2}(x,Q^{2})$ and the polarized structure function $g_{1}(x,Q^{2})$ for the first time. These results are relevant for precision QCD analyses of the World deep-inelastic data and the data taken at future colliders, such as the Electron–Ion Collider, since the scaling violations due to massless and massive Wilson coefficients are significantly different. In order to measure the strong coupling constant $\alpha_{s}(M_{Z}^{2})$ and the twist-2 parton distribution functions consistently at highest precision, the next-to-next-to-leading order corrections have to be taken into account. Furthermore, the complete three-loop corrections will allow to reduce the present theory error of the charm mass $m_{c}$ as measured form deep-inelastic data. We provide a fast and precise public numerical code for the unpolarized and polarized massive Wilson coefficients in the asymptotic region.
1 Introduction
@@
With the final results on the next-to-next-to-leading order (NNLO) single-mass corrections to the heavy-flavor Wilson coefficients in the region $Q^{2}\gg m_{Q}^{2}$, where $m_{Q}$ is the heavy-quark mass, numerical predictions on the structure functions $F_{2}$ and $g_{1}$ can be made at this order for the first time.
As has been shown in Ref. [5], the extraction of the charm quark mass from DIS data relying on approximate representations [9] leads to a large theory uncertainty of $\delta m_{c}^{\rm theory}=\tiny{\begin{array}[]{c}+0.00\\ -0.07\end{array}}\penalty 10000\ \rm GeV$ while the experimental error is $0.03\penalty 10000\ \rm GeV$. The experimental error will even decrease with new data from the Electron-Ion Collider (EIC) in the future [10, 11]. Moreover, a series of Wilson coefficients, which are numerically smaller were not considered in this analysis. The completed NNLO calculation will lower this uncertainty.
The purpose of this paper is to compile the single-mass three-loop contributions to the heavy-flavor Wilson coefficients to quantify the effect of the different contributions prior to analyses of experimental data. In the polarized case our results prepare for future precision measurements at the EIC. These quantifications were not given before. Furthermore, a public numerical library for the massive Wilson coefficients is provided.
The massive Wilson coefficients can be represented by Mellin convolutions of the massive operator matrix elements (OMEs) and pieces of the massless Wilson coefficients. The asymptotic single-mass Wilson coefficients are given by, cf. Ref. [12],
$L_{2,q}^{\rm NS}$ $=$ $a_{s}^{2}\left[A_{qq,Q}^{{\rm NS},(2)}+\hat{C}^{{\sf NS},(2)}_{2,q}\right]+a_{s}^{3}\left[A_{qq,Q}^{{\rm NS},(3)}+A_{qq,Q}^{{\rm NS},(2)}C_{2,q}^{{\rm NS},(1)}+\hat{C}^{{\rm NS},(3)}_{2,q}\right$], (1)
${\tilde{L}}_{2,q}^{\rm PS}$ $=$ $a_{s}^{3}\left[\penalty 10000\ \tilde{A}_{qq,Q}^{{\rm PS},(3)}+A_{gq,Q}^{(2)}\penalty 10000\ \penalty 10000\ \tilde{C}_{2,g}^{(1)}+\hat{\tilde{C}}^{{\rm PS},(3)}_{2,q}\right$], (2)
${\tilde{L}}_{2,g}$ $=$ $a_{s}^{2}A_{gg,Q}^{(1)}\tilde{C}_{2,g}^{(1)}+a_{s}^{3}\Bigl[\penalty 10000\ \tilde{A}_{qg,Q}^{(3)}+A_{gg,Q}^{(1)}\penalty 10000\ \penalty 10000\ \tilde{C}_{2,g}^{(2)}+A_{gg,Q}^{(2)}\penalty 10000\ \penalty 10000\ \tilde{C}_{2,g}^{(1)}+\penalty 10000\ A_{Qg}^{(1)}\penalty 10000\ \penalty 10000\ \tilde{C}_{2,q}^{{\rm PS},(2)}+\hat{\tilde{C}}^{(3)}_{2,g}\Bigr$], (3)
$H_{2,q}^{\rm PS}$ $=$ $a_{s}^{2}\left[\penalty 10000\ A_{Qq}^{{\sf PS},(2)}+\penalty 10000\ \tilde{C}_{2,q}^{{\sf PS},(2)}\right]+a_{s}^{3}\Bigl[\penalty 10000\ A_{Qq}^{{\rm PS},(3)}+\penalty 10000\ \tilde{C}_{2,q}^{{\rm PS},(3)}+A_{gq,Q}^{(2)}\penalty 10000\ \tilde{C}_{2,g}^{(1)}+A_{Qq}^{{\rm PS},(2)}\penalty 10000\ C_{2,q}^{{\sf NS},(1)}\Bigr],$ (4)
$H_{2,g}$ $=$ $a_{s}\left[A_{Qg}^{(1)}+\tilde{C}_{2,g}\right]+a_{s}^{2}\left[A_{Qg}^{(2)}+A_{Qg}^{(1)}C_{2,q}^{{\rm NS},(1)}+A_{gg,Q}^{(1)}\tilde{C}_{2,g}^{(1)}+\tilde{C}_{2,g}^{(2)}\right]+a_{s}^{3}\left[A_{Qg}^{(3)}+A_{Qg}^{(2)}C_{2,q}^{{\rm NS},(1)}\right$. (5)
$+A_{gg,Q}^{(2)}\tilde{C}_{2,g}^{(1)}+A_{Qg}^{(1)}\left(C_{2,q}^{{\rm NS},(2)}+\tilde{C}_{2,q}^{{\rm PS},(2)}\right)\left.+A_{gg,Q}^{(1)}\tilde{C}_{2,g}^{(2)}+\tilde{C}_{2,g}^{(3)}\right]$
in Mellin space, where the Mellin convolutions simplify to ordinary products. Here $A_{ij}^{(k)}$ denote the expansion coefficients of the massive OME $\langle j|O^{i}|j\rangle$ of a local twist-2 operator $O$ located at the line $i=Q,q,g$ with the external state $j=q,g$, where $Q$ denotes a massive quark, $q$ a massless quark, $g$ the gluon. The expansion coefficients of the massless Wilson coefficients $C_{i}^{l,(k)}$ depend on $N_{F}+1$ flavors and $\tilde{f}(N_{F})=f(N_{F})/N_{F}$. For the other massive Wilson coefficients see Refs. [13, 12]. The massive OMEs have been calculated to three-loop order in the unpolarized cases in Refs. [14, 19, 15, 16, 9, 17, 18] and in the polarized cases in Refs. [19, 21, 20, 17, 18]. [^1] The one– and two–loop corrections were calculated in Refs. [22, 23, 24, 25, 26, 27] and [30, 31, 13, 28, 29, 32, 33, 34].
@@
Beyond the massive OMEs contributing to the massive Wilson coefficients, there are also others, $(\Delta)A_{gq,Q}$ and $(\Delta)A_{gg,Q}$, playing a role in the variable flavor number scheme [35, 16, 36, 38, 39, 37].
@@
It has been shown in Ref. [13] that for the structure function $F_{2}(x,Q^{2})$ the heavy-flavor corrections in the asymptotic region $Q^{2}\gg m_{Q}^{2}$ agree with the complete result at the level of $O(1\%)$ if $Q^{2}/m_{Q}^{2}>10$ at two-loop order. This cut also safely removes the higher-twist effects, cf. Refs. [40, 41, 42]. The structure function $F_{2}(x,Q^{2})$ consists of the massless part, $F_{2}^{\rm light}(x,Q^{2})$ for the light flavors, and the massive part, $F_{2}^{\rm heavy}(x,Q^{2})$, where the Wilson coefficients have either a charm or a bottom quark contribution. [^2] Here we calculate the ‘extrinsic’ heavy-flavor contributions, but not intrinsic heavy-flavor corrections, as suggested by a Fock-space formalism in the infinite-momentum frame [43, 44].
@@
In our illustrations, we work in the fixed flavor number scheme for $N_{F}=3$ light flavors with charges $e_{k}$. The NkLO contribution to $F_{2}^{\text{total}}$ combines $F_{2}^{\text{light}}$ contributions through $O(a_{s}^{k})$ and $F_{2}^{\text{heavy}}$ contributions through $O(a_{s}^{k+1})$ for $k=0,1,2$. The formal structures are the same for the structure functions $F_{2}(x,Q^{2})$ and $g_{1}(x,Q^{2})$, and are obtained by substituting the unpolarized building blocks with the polarized ones. For $g_{1}(x,Q^{2})$ the prefactor $x$ in Eqs. (7,10) has to be replaced by $1/2$.
The light-flavor part of $F_{2}$ is given by [45, 46]
$F_{2}^{\rm light}(x,Q^{2})$ $=$ $x\sum_{k=1}^{N_{F}}e_{k}^{2}\Biggl\{\frac{1}{N_{F}}\Biggl[C_{2,q}^{\rm S}\left(x,\frac{Q^{2}}{\mu^{2}}\right)\otimes\Sigma(x,\mu^{2})+C_{2,g}^{\rm S}\left(x,\frac{Q^{2}}{\mu^{2}}\right)\otimes G(x,\mu^{2})\Biggr]$ (7)
@@
$\Delta_{k}(x,\mu^{2})$ $=$ $f_{k}(x,\mu^{2})+{f}_{\bar{k}}(x,\mu^{2})-\frac{1}{N_{F}}\Sigma(x,\mu^{2}),$ (8)
$\Sigma(x,\mu^{2})$ $=$ $\sum_{k=1}^{N_{F}}[f_{k}(x,\mu^{2})+{f}_{\bar{k}}(x,\mu^{2})],$ (9)
with $f_{k,(\bar{k})}$ the quark and antiquark number densities, and $\mu$ denotes the factorization scale, which we set equal to the renormalization scale. The unpolarized massless Wilson coefficients $C_{2,i}^{j}$ were calculated to three-loop order in Refs. [45, 46] and the polarized ones $\Delta C_{1,i}^{j}$ in Ref. [46]. The heavy-flavor part is given by, cf. Ref. [12],
$F_{2}^{Q}(x,Q^{2})=$ (10)
$x\Biggl\{\sum_{k=1}^{N_{F}}e_{k}^{2}\Biggl[L_{2,q}^{\rm NS}(x)\otimes(f_{k}(x,\mu^{2})+f_{\bar{k}}(x,\mu^{2}))+\frac{1}{N_{F}}\left({L}_{2,q}^{\rm PS}(x)\otimes\Sigma(x,\mu^{2})+{L}_{2,g}^{\rm S}(x)\otimes G(x,\mu^{2})\right)\Biggr]$
@@
It contains both flavor non-singlet and singlet contributions. [^3] For the study of pure non-singlet structure functions, see Ref. [47]. There the heavy-flavor corrections are small. Here we have suppressed the dependence on $Q^{2},m_{Q}^{2}$ and $\mu^{2}$ in the heavy-flavor Wilson coefficients $H_{i}$ and $L_{i}$, and $\otimes$ denotes the Mellin convolution, [^4] For the evaluation of the convolution integrals in the numerical illustrations we use the integrator of Ref. [48].
$A(x)\otimes B(x)=\int_{0}^{1}dx_{1}\int_{0}^{1}dx_{2}\delta(x-x_{1}x_{2})A(x_{1})B(x_{2}).$ (11)
The following convolutions contribute, cf. Eqs. (256–258) of Ref. [34],
$\left(\left[\frac{f}{1-x}\right]_{+}\otimes g\right)(x)$ $=$ $\int_{x}^{1}dz\frac{f(z)}{1-z}\left[\frac{1}{z}g\left(\frac{x}{z}\right)-g(x)\right]-g(x)\int_{0}^{x}dz\frac{f(z)}{1-z},$ (12)
$(h\otimes g)(x)$ $=$ $\int_{x}^{1}\frac{dz}{z}h(z)g\left(\frac{x}{z}\right),$ (13)
$(\delta(1-x)\otimes g)(x)$ $=$ $g(x).$ (14)
Here $f(x)$ and $g(x)$ are functions with support $x\in]0,1[$.
@@
In Mellin space, the massive Wilson coefficients are polynomials of the expansion coefficients of the massive OMEs $(\Delta)A_{ij}$ and expansion coefficients of the massless Wilson coefficients, see Refs. [12, 16]. The massless Wilson coefficients to three-loop order can all be expressed in terms of either harmonic sums [49, 50] in Mellin space or harmonic polylogarithms [51] in $x$-space. For the massive Wilson coefficients, this only applies to the $N_{F}$-contributions [14] and to $(\Delta)L_{q}^{\rm NS},(\Delta)L_{q}^{\rm PS}$ and $(\Delta)L_{g}^{\rm S}$. The pure-singlet Wilson coefficients $(\Delta)H_{q}^{\rm PS}$ depend on generalized harmonic sums [52, 53] in Mellin space and can be cast into harmonic polylogarithms with changed argument in $x$-space. The gluonic Wilson coefficients $(\Delta)H_{g}^{\rm S}$ receive also finite binomial sum [54] and higher transcendental function contributions in Mellin space. They contain square-root valued iterated integrals [54] and $_{2}F_{1}-$solutions [55] in $x$-space. [^5] For surveys on the different mathematical structures and the used computer algebraic methods see Refs. [56, 57]. Since we calculate the heavy-flavor corrections to inclusive structure functions, there are also contributions with massless final states, but virtual heavy-flavor corrections.
The paper is organized as follows. In Section 2 we compute the single-mass heavy-quark corrections for the unpolarized structure function $F_{2}(x,Q^{2})$ and in Section 3 those for the polarized structure function $g_{1}(x,Q^{2})$ to three-loop order. The details of the fast and precise numerical code provided, which is well suited for QCD-analyses of hard-scattering data, are described in Section 4 and the conclusions are contained in Section 5.
2 The Structure Function $F_{2}(x,Q^{2})$
In the unpolarized case, we use the parton distribution functions (PDFs) of Ref. [7] for $N_{F}=3$. [^6] From LHAPDF [58].
Figure 2: The different heavy-flavor contributions due to charm quarks to the structure function $F_{2}(x,Q^{2})$ by the Wilson coefficients $H_{g}^{\rm S},H_{q}^{\rm PS},L_{q}^{\rm NS},L_{g}^{\rm S}$ and $L_{g}^{\rm PS}$ at $Q^{2}=100\penalty 10000\ \rm GeV^{2}$ at different orders in the strong coupling constant up to $O(a_{s}),O(a_{s}^{2})$ and $O(a_{s}^{3})$.
For the values of $\alpha_{s}$ in the Wilson coefficients, we use flavor matching with the values $\alpha_{s}(25\penalty 10000\ \rm GeV^{2})=0.2020,\alpha_{s}(10^{2}\penalty 10000\ \rm GeV^{2})=0.1706$, and $\alpha_{s}(10^{3}\penalty 10000\ \rm GeV^{2})=0.1359$, corresponding to $\alpha_{s}(M_{Z}^{2})=0.1147$, consistent with the PDF fits. The massive Wilson coefficients were calculated referring to on-shell heavy-quark masses. We use the following values
$m_{c}=1.59\penalty 10000\ \rm GeV,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ {\it m_{b}}=4.78\penalty 10000\ \rm GeV,$ (15)
@@ 2 The Structure Function $F_{2}(x,Q^{2})$
In Figure 1, we illustrate the prediction for the total structure function $F_{2}^{\rm total}(x,Q^{2})$, including the single-mass charm and bottom quark contributions. In the region $x\sim 10^{-5}$, it rises from $Q^{2}=25\penalty 10000\ \rm GeV^{2}$ to $Q^{2}=1000\penalty 10000\ \rm GeV^{2}$ from values of $\sim 3.5$ to 10.5.
Figure 2 shows the contributions of the different heavy-flavor Wilson coefficients to $F_{2}^{c}$ at $Q^{2}=100\penalty 10000\ \rm GeV^{2}$. The different lines illustrate also the contributions from $O(a_{s})$ to $O(a_{s}^{3})$ to the structure function. The largest contribution is due to the gluonic Wilson coefficient $H_{g}^{\rm S}$ which was scaled down by a factor of 20 for better readability. The corrections are positive and grow with the order in $a_{s}$. It is followed in size by a negative contribution from the pure-singlet Wilson coefficient $H_{q}^{\rm PS}$, which starts at $O(a_{s}^{2})$. In the large-$x$ region, the non-singlet Wilson coefficient $L_{q}^{\rm NS}$ yields the dominant part.
@@
Around $x=0.1$ there is a turning point above which the structure function takes lower values for larger values of $Q^{2}$. However, the corrections compared to $F_{2}^{c,b}$ stay rather small. The corrections can become negative since virtual effects of heavy quarks with massless final states are contained in $F_{2}^{c}$ and $F_{2}^{b}$ and some Wilson coefficients start at $O(a_{s}^{2})$ only. Also the contributions of $L_{g}^{\rm S}$ and $L_{g}^{\rm PS}$ are of importance at the level of accuracies of $O(1\%)$. $L_{g}^{\rm S}$ is tiny at $O(a_{s}^{2})$, but receives a relative large correction at $O(a_{s}^{3})$. A part of Wilson coefficients emerging at higher order in the coupling constant $a_{s}$ can be negative as long as the unpolarized structure function remains positive.
In Figure 3 the heavy-flavor corrections to $F_{2}(x,Q^{2})$ are shown at NNLO. At $x=10^{-5}$, they grow from 1 to $\sim 4$ for $Q^{2}=25\penalty 10000\ \rm GeV^{2}$ to $Q^{2}=1000\penalty 10000\ \rm GeV^{2}$. In the large-$x$ region the corrections are negative.
@@
at NNLO. In the small-$x$ region of $x\sim 10^{-5}$, the fractions of the heavy-flavor corrections evolve from $Q^{2}=25\penalty 10000\ \rm GeV^{2}$ to $Q^{2}=1000\penalty 10000\ \rm GeV^{2}$ from 26 % to 41 %. At $x\sim 0.2$, the heavy-flavor corrections become zero, which is due to a combination of real and virtual corrections at NLO and NNLO. In the very-large-$x$ region, the heavy-flavor contributions at NNLO become negative and as large as $10\%$, but the structure function $F_{2}$ itself remains non-negative. It is clearly seen that the scaling violations of the massive contributions are rather different compared to those in the massless case up to high values of $Q^{2}$.
3 The Structure Function $g_{1}(x,Q^{2})$
In the polarized case the calculation is performed in the Larin scheme [60]. Both for the massive three-loop OMEs and the massless three-loop Wilson coefficients, the scheme transformations to the $\overline{\rm MS}$ scheme are not yet known. However, if all contributing quantities, including also the